# https://www.dekelgalor.com/noise2image

### v1

@@ 5.1 Results
Noise2Image, synthetic 20.9 0.589 0.462 16.3 0.468 0.583
@@
Table 1: Quantitative results for static scene reconstruction with in-distribution and out-of-distribution testing data. Our Noise2Image models are trained with either synthetic or experimental noise data. Pre-trained events-to-video reconstruction (E2VID) methods [^3] Since E2VID models were trained by synthetic motion data generated from the MS-COCO dataset [32], both of our testing sets are considered as out-of-distribution.are used as baselines. We report peak signal-to-noise ratio (PSNR), structured similarity (SSIM), and perceptual similarity (LPIPS [54]).
@@
Figure 4: Comparison between Noise2Image and baseline pre-trained event-to-video (E2VID) methods on noise event-to-intensity reconstruction. The first row shows the input event count (captured in experiment) aggregated over a 1-second window. Noise2Image is trained using either synthetic or experimental data as specified. Full comparison with all baseline methods can be found in Supplementary Fig. 2.
@@
For static-only scenes, the quantitative scores for Noise2Image as well as baseline methods using NE2I dataset are reported in Table 3. As shown in Fig. 4, the Noise2Image model trained by either experimental or synthetic data recovers detailed intensity images from the count of noise events. Both Noise2Image models generalized well to out-of-distribution testing data, providing a good level of contrast and details, suggesting that Noise2Image learns the correlation between noise event and intensity. The Noise2Image model trained by experimental data out-performed the one trained by the synthetic data by 4.1dB of PSNR. As the synthetic data-trained model often recovered speckle-like patterns for the uniform image background, we speculate that this performance margin is caused by the spatial correlation of the noise events, which is beyond our synthesis model but potentially captured by the experimental data training. Baseline E2VID methods did not perform well on static scene reconstruction for two reasons: they were only trained by scenes with various degrees of motion (not static scenes) and the training data of E2VID was generated without an intensity-dependent noise model [41].
@@
In addition to testing data, we tested Noise2Image on real-world scenes, both indoor and outdoor. The Noise2Image model trained by images displayed on a monitor generalizes well for scenes outside of the laboratory setting, as shown in Fig. 1 and Supplementary Fig. 3. We have implemented a real-time demo running on a laptop computer as well.
5.1.1 Effect of aggregation duration.
@@
Table 2 shows the effect of varying the aggregation window time for training and testing the Noise2Image method. As the aggregation duration shortens, fewer noise events will be triggered, and the estimation of the true event count becomes less accurate. Table 2 shows that the Noise2Image reconstruction works even with short aggregation duration (0.1 seconds), although longer integration will result in better reconstruction quality.
@@
Table 2: Noise2Image performance using noise event counts aggregated over different time windows. For each aggregation duration, the Noise2Image model is trained and evaluated using the experimental data with identical training setting.
5.1.2 Reconstructing scenes with dynamic components.
Finally, we demonstrated that Noise2Image is complementary to E2VID reconstruction when the scene has both static and dynamic components. We imaged a fan rapidly moving in front of a static scene, and fed signal events into a pre-trained E2VID model [42]. While the E2VID method performed well on the moving object, it could not recover the background static scene (Fig. 5c). To incorporate Noise2Image, we first identified a motion mask at each timepoint (Fig. 5b) by thresholding signal events. We then aggregated noise events and fed the event count at pixels without motion to the Noise2Image model. Stitching together the moving foreground from E2VID and the static background from Noise2Image, we obtained the final high-quality reconstruction as in Fig. 5d and Supplementary Video 1.
Figure 5: Dynamic scene reconstruction of a moving fan in front of a static scene (see Supplementary Video 1). a. Aggregated event count. The first row is at a timepoint with no motion and only faint noise, and the second row has motion. b. Motion mask obtained from thresholding the signal events as determined by the event denoiser. c. E2VID reconstruction using pre-trained model from [42]. d. The dynamic foreground reconstructed by E2VID and the static background reconstructed by Noise2Image are stitched together.
6 Discussion
Denoising and camera bias parameters [35] can also affect this illuminance dependency. Our data was acquired without denoising and using bias parameters default to our event camera. We also tested different high-pass filtering bias values which affect the correlation between noise event rate and illuminance (Supplementary Fig.4). For this reason, the trained Noise2Image model is bound to the camera model and bias parameters that were used in the training data collection. We hope that our modeling of noise events can be improved upon by incorporating the analog nature of event sampling and modeling additional event camera bias parameters, e.g. using non-parametric models to account for unknown camera behavior.
One practical constraint for our study is that the dynamic range of Noise2Image is limited by the monitor we used. As a result, the increase in noise events at high illuminance settings due to leakage current cannot be measured using our setup. This issue can likely be alleviated by changing to a high-dynamic range and high-brightness monitor or projector for data collection. We hypothesize that our Noise2Image approach will still be valid for brighter scenes since the statistics of leakage current-induced noise events is quite distinct [36].
Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1752814. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Research was supported by the National Eye Institute of the National Institute of Health under award number R00EY032179. Ruiming Cao was supported in part by Siebel Scholarship. This work was supported by Weill Neurohub Investigators Program. This work has been made possible in part by CZI grant DAF2021-225666 and grant DOI 10.37921/192752jrgbnh from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation (funder DOI 10.13039/100014989). Laura Waller is a Chan Zuckerberg Biohub SF Investigator.
Supplementary Fig. 3: Real-world examples of Noise2Image taken outside of the laboratory setting. The Noise2Image model trained by synthetic data results in background artifacts on the in-door scene (second row, second column), presumably caused by the one-to-many relationship of Eq. 3. In contrast, the Noise2Image trained by experimental data predicts the background correctly, hinting that there exist spatial correlations in the experimental noise events, beyond our derived spatially independent noise event synthesis. The reference images were taken by an iPhone 12 plus back camera.

### v2

@@ 5.2 Results
Noise2Image, synthetic 20.3 0.614 0.449 16.6 0.504 0.532
@@
Table 1: Quantitative results for static scene reconstruction with in-distribution and out-of-distribution testing data. Our Noise2Image models are trained with either synthetic or experimental noise data. Pre-trained events-to-video reconstruction (E2VID) methods are used as baselines. We report peak signal-to-noise ratio (PSNR), structured similarity (SSIM), and perceptual similarity (LPIPS [57]). We note that since E2VID models were trained by synthetic motion data generated from the MS-COCO dataset [58], both of our testing sets are considered out-of-distribution.
@@
Figure 4: Comparison between Noise2Image and baseline pre-trained event-to-video (E2VID) methods on noise event-to-intensity reconstruction. The first row shows the input event count (captured in experiment) aggregated over a 1-second window. Noise2Image is trained using either synthetic or experimental data as specified. Full comparison with all baseline methods can be found in Fig. S2.
@@
For static-only scenes, the quantitative scores for Noise2Image as well as baseline methods using NE2I dataset are reported in Table 1. As shown in Fig. 4, the Noise2Image model trained by either experimental or synthetic data recovers detailed intensity images from the count of noise events. Both Noise2Image models generalized well to out-of-distribution testing data, providing a good level of contrast and details, suggesting that Noise2Image learns the correlation between noise event and intensity. The Noise2Image model trained by experimental data out-performed the one trained by the synthetic data by 4.1dB of PSNR. As the synthetic data-trained model often recovered speckle-like patterns for the uniform image background, we speculate that this performance margin is caused by the variability of contrast thresholds across pixels, which is beyond our synthesis model but potentially captured by the experimental data training. Baseline E2VID methods did not perform well on static scene reconstruction for two reasons: they were only trained by scenes with various degrees of motion (not static scenes) and the training data of E2VID was generated without an intensity-dependent noise model [32].
Figure 5: Real-world examples of Noise2Image taken outside of the laboratory setting. The Noise2Image model trained by synthetic data results in background artifacts on the in-door scene (second row, second column), presumably caused by the one-to-many relationship of Eq. 3. In contrast, the Noise2Image trained by experimental data predicts the background correctly, hinting that there exist spatial structures in the experimental noise sensitivity, beyond our derived spatially independent noise event synthesis. The reference images were taken by an iPhone 12 plus back camera.
@@
In addition to testing data, we tested Noise2Image on real-world scenes, both indoor and outdoor. The Noise2Image model trained by images displayed on a monitor generalizes well for scenes outside of the laboratory setting, as shown in Fig. 5. We have implemented a real-time demo running on a laptop computer as well.
@@
Table S1 shows the effect of varying the aggregation window time for training and testing the Noise2Image method. As the aggregation duration shortens, fewer noise events will be triggered, and the estimation of the true event count becomes less accurate. Noise2Image reconstruction works even with short aggregation duration (0.1 seconds), although longer integration will result in better reconstruction quality. We also tested using only events from a single polarity, which resulted in a slight reduction in performance as shown in Table 2.
Event polarity In-distribution test Out-of-distribution test
PSNR SSIM LPIPS PSNR SSIM LPIPS
Positive & negative 25.0 0.742 0.349 19.3 0.552 0.509
Positive only 24.1 0.686 0.378 18.7 0.501 0.526
Negative only 24.2 0.696 0.378 18.8 0.513 0.531
Table 2: Ablation study of event polarities. We train and test Noise2Image method using both positive and negative events, positive events only, and negative events only.
5.3 Reconstructing scenes with dynamic components
Finally, we demonstrated that Noise2Image is complementary to E2VID reconstruction when the scene has both static and dynamic components. We imaged a fan rapidly moving in front of a static scene, and fed signal events into a pre-trained E2VID model [25]. While the E2VID method performed well on the moving object, it could not recover the background static scene (Fig. 6c). To incorporate Noise2Image, we first identified a motion mask at each timepoint (Fig. 6b) by thresholding signal events. We then aggregated noise events and fed the event count at pixels without motion to the Noise2Image model. Stitching together the moving foreground from E2VID and the static background from Noise2Image, we obtained the final high-quality reconstruction as in Fig. 6d and Visualization 1.
Figure 6: Dynamic scene reconstruction of a moving fan in front of a static scene (see Visualization 1). a. Aggregated event count. The first row is at a timepoint with no motion and only faint noise, and the second row has motion. b. Motion mask obtained from thresholding the signal events as determined by the event denoiser (YNoise denoiser from [59] implemented in [35]). c. E2VID reconstruction using pre-trained model from [25]. d. The dynamic foreground reconstructed by E2VID and the static background reconstructed by Noise2Image are stitched together.
6 Discussion
Denoising and camera bias parameters [60] can also affect this illuminance dependency. Our data was acquired without denoising and using bias parameters default to our event camera. We also tested different high-pass filtering bias values which affect the correlation between noise event rate and illuminance (Fig. S3). For this reason, the trained Noise2Image model is bound to the camera model and bias parameters that were used in the training data collection. We hope that our modeling of noise events can be improved upon by incorporating the analog nature of event sampling and modeling additional event camera bias parameters, e.g., using non-parametric models to account for unknown camera behavior.
In our evaluation, Noise2Image trained by experimental data outperforms the one trained by synthetic data, which suggests discrepancy between the captured noise distribution and our theoretical noise model. From our observation, there exist some spatial structures in the experimental noise event count which is beyond our theoretical noise model. We believe this is caused by the variability of contrast threshold, a known issue related to the manufacturing of event sensor [34, 61].
The dynamic range of Noise2Image is limited to the moderate-brightness regime, where noise events are abundant. Although our theoretical noise model still holds in the high-brightness regime, there are fewer noise events, providing little signal for static scene recovery. Our empirical study is further constrained by the dynamic range of the monitor we used, suggesting that a high-dynamic-range, high-brightness monitor or projector could enhance the dynamic range of the experimental training data. We hypothesize that our Noise2Image approach may still work for brighter scenes with the presence of leakage current-induced noise events [14], and will leave this for future research. For low-brightness scenes, it is worth further investigating the feasibility of the Noise2Image approach under extremely low-signal conditions, such as in fluorescence microscopy.
Disclosure
Authors declare no conflict of interest.
Code and Data Availability
Code and data presented in this paper are available: https://github.com/rmcao/noise2image.
Acknowledgments
We thank Joshua Follansbee for the feedback on our derivations. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1752814. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Research was supported by the National Eye Institute of the National Institute of Health under award number R00EY032179. Ruiming Cao was supported in part by Siebel Scholarship. This work was supported by Weill Neurohub Investigators Program. This work has been made possible in part by CZI grant DAF2021-225666 and grant DOI 10.37921/192752jrgbnh from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation (funder DOI 10.13039/100014989). Laura Waller is a Chan Zuckerberg Biohub SF Investigator.
@@ Acknowledgments
Supplementary Table 1: Noise2Image performance using noise event counts aggregated over different time windows. For each aggregation duration, the Noise2Image model is trained and evaluated using the experimental data with identical training setting.